# GOALS-JWST: The Warm Molecular Outflows of the Merging Starburst Galaxy NGC 3256

### v1

@@ 5.3 Gas Conditions in the Outflow
To investigate the impact the outflows have on the local ISM, we plot [Fe ii]/Pa versus H/Br line ratios in Figure 7. Higher [Fe ii] and H line ratios relative to the recombination lines indicate stronger gas excitation, and are located at the southern edge of the FoV, where they are spatially coincident with the blueshifted outflow. This region is also where the outflow flux is the strongest (see Figure 4) and has the highest fraction of warm gas (see Figure 6). This increase in the H and [Fe ii] line ratios could be caused by the outflow impacting a denser region in the ISM, possibly in an interaction with the disk of the northern galaxy. This interpretation can be supported by the comparatively low value of the power-law index of the southern outflow, 4.54, which is lower than the mean found in other galaxies, 4.84 (Togi & Smith, 2016). Indices in this range are seen in other cases where shocks are present (Appleton et al., 2017), suggesting a shock front created by the outflow. However, most of the outflow falls outside the NIRSpec FoV, so it is difficult to assess the full effect of the excitation using this method. While the narrow FoV limits us from analyzing the redshifted outflow as well, the outflow regions that do fall within the NIRSpec FoV of the north galaxy also show elevated values for these line ratios, indicating that the northwards outflow may also be energizing the ISM.
@@ 5.4.1 Energy of the Outflow
Setting , 90 km s, and 750 km s gives a combined outflow energy of erg. For the cold H outflow, where , and assuming the same kinematics yields erg. Assuming the outflow time scale to be yr, this results in a kinetic power of erg s and erg s.
Put together, the temperatures, kinematics, and energetics of the outflows can provide clues to their origin. As shown in Figure 6, we find the highest warm gas fractions close to the S nucleus, with fractions decreasing at larger radii, an indication of heating from the nuclear source. Moreover, in addition to the symmetry to the north and south of the S nucleus, the fast velocities (10 km s), large mass loading factor (), and power ( erg s) found here are all typical of AGN-driven outflows (Rodríguez-Ardila et al., 2006; Müller-Sánchez et al., 2011; Hill & Zakamska, 2014; Bae & Woo, 2016; Harrison et al., 2016; Veilleux et al., 2020).
@@ 5.4.2 Feedback on Star Formation
It is well known that sustained outflows can suppress star formation (e.g., King & Pounds, 2015; Armus et al., 2020). A commonly used IR tracer of star formation are PAHs (Peeters et al., 2004; Inami et al., 2018; Lai et al., 2020, 2022). Because PAH grains are fragile, they can easily be destroyed in harsh radiation environments. As such, we can utilize PAH fluxes to assess if the outflows have any significant feedback on the local ISM. For this, we opt to use the 3.3 and 6.2 m PAHs since they are isolated from other features and offer the best spatial resolution among the PAHs available in JWST.
@@
In the right panel of Figure 7, we plot 3.3 m PAH flux contours over the excitation map. Most of the PAH flux is centered around the S nucleus with emission extending to the east and west, likely tracing the disk of the southern galaxy. The emission seems to decrease in the outflow regions but the limited FoV makes it difficult to ascertain the exact correlation between the outflow and PAH emission. As such, we can take advantage of a larger FoV of MIRI to examine the PAH at 6.2 m, as shown in the flux map of Figure 8. Like PAH 3.3 m, we measure the strongest PAH emission at the nuclear region of each galaxy, with other prominent emission seen in the disks and spiral arms, likely caused by the star formation occurring there. As was done in Figure 4, we overlay the outflowing gas as traced by the H S(1) 17.0 m. While the PAH emission is generally lower in the regions coincident with the outflow, we see no clear disruption in the overall PAH structure. The PAH emission clearly traces the disks and spiral arms, and regions with lower emission are likely due to the geometry of the system. Although the outflows we detect are fast and significantly heat the molecular gas, we do not see any clear evidence of them inducing negative feedback on the local star formation.
@@ 6 Conclusion
Utilizing the exceptional resolution and sensitivity of JWST, we investigate the nearby merging LIRG NGC 3256 on scales of 100 pc. Here, we showcase the capability of JWST to spatially resolve the MIR H lines, allowing us to directly identify and trace the outflowing warm H gas. Using the H lines from 5.0–17.0 m, we analyze the mass, temperature distribution, kinematics, and energetics of the outflowing warm H gas, and assess the impact of the outflows on the local star formation. The following is a summary of our findings:
• Warm, outflowing H gas is detected in a collimated outflow originating from the southern nucleus. The outflows extend out to a distance of 700 pc, and have a deprojected maximum velocity of 1,000 km s. The emission is most intense near the far edges of the outflows, possibly due to impacting the disk of the northern galaxy. We do not detect any significant outflowing H gas originating from the northern nucleus.
• Using the extensive wavelength coverage of JWST, we use the full set of S(8)–S(1) H lines to directly calculate the outflowing warm H gas mass: = 8.910, where the / mass fraction is 4. With an outflow time scale of about yr, the outflow mass rate is about = 1.3 M yr. These masses yield an outflow energy of erg and kinetic power of erg s. The mass loading factor of the warm outflowing gas is roughly unity. However, when taking into account the cold H component, it could 5 or higher.
• The spatial map of the outflowing H S(7)/S(1) flux ratio reveals higher ratios closer to the S nucleus. This indicates a larger fraction of warmer gas is closer to the nucleus and decreases as the outflow expands outwards, suggesting that the S nucleus is the heating source of the outflowing gas.
• Analysis of the gas excitation in the outflow regions using [Fe ii]/Pa and H/Br line ratios shows enhanced ratios compared to those seen in typical star forming regions, indicating the outflow is energizing these regions, possibly as a shock. The value of the power-law index is also indicative of shocks, particularly in the blueshifted outflow. However, flux maps of the 6.2 m PAH do not show the outflowing gas having a clear negative feedback effect on the local star formation.
@@
The JWST data presented in this paper were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program 1328 and are supported by NASA grant ERS-01328. They can be accessed at 10.17909/8mkr-xr82 (catalog https://doi.org/10.17909/8mkr-xr82). STScI is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS5–26555. Support to MAST for these data is provided by the NASA Office of Space Science via grant NAG5–7584 and by other grants and contracts. T.B. and H.I. acknowledge support from JSPS KAKENHI Grant Number JP21H01129 and the Ito Foundation for Promotion of Science. A.S.E. and S.T.L. acknowledge support from NASA grants HST-GO15472 and HST-GO16914. V.U. acknowledges partial funding support from NASA Astrophysics Data Analysis Program (ADAP) grant Nos. 80NSSC20K0450 and 80NSSC23K0750 and STScI grant Nos. HST-AR-17063.005-A, HST-GO-17285.001-A, and JWST-GO-01717.001-A. C.R. acknowledges support from Fondecyt Regular grant 1230345 and ANID BASAL project FB210003 The Flatiron Institute is supported by the Simons Foundation. S.A. gratefully acknowledges support from an ERC Advanced Grant 789410, the Swedish Research Council and the Knut and Alice Wallenberg (KAW) foundation. F.M-S. acknowledges support from NASA through ADAP award 80NSSC19K1096. This paper makes use of the following ALMA data: ADS/JAO.ALMA2011.0.00525.S and 2015.1.00902.S. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), NSTC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ.

### v3

@@ 5.3 Gas Conditions in the Outflow
To investigate the impact the outflows have on the local ISM, we plot [Fe II]/Pa versus H2/Br line ratios in Figure 7. Higher [Fe II] and H2 ratios relative to the recombination lines indicate stronger gas excitation, and are located at the southern edge of the FoV, where they are spatially coincident with the blueshifted outflow. This region is also where the outflow flux is the strongest (see Figure 4) and has the highest fraction of warm gas (see Figure 6). This increase in the H2 and [Fe II] ratios could be caused by the outflow impacting a denser region in the ISM, possibly in an interaction with the disk of the northern galaxy. This interpretation can be supported by the comparatively low value of the power-law index of the southern outflow, 4.54, which is lower than the mean found in other galaxies, 4.84 (Togi & Smith, 2016). Indices in this range are seen in other cases where shocks are present (Appleton et al., 2017), suggesting a shock front created by the outflow. However, most of the outflow falls outside the NIRSpec FOV, so it is difficult to assess the full effect of the excitation using this method. While the narrow FOV limits us from analyzing the redshifted outflow as well, the outflow regions that do fall within the NIRSpec FOV of the north galaxy also show elevated values for these line ratios, indicating that the northwards outflow may also be energizing the ISM.
@@ 5.4.1 Energy of the Outflow
Setting , 90 km s-1, and 730 km s-1 gives a combined outflow energy of erg. For the cold H2 outflow, where , and assuming the same kinematics yields erg. Assuming the outflow time scale to be yr, this results in a kinetic power of erg s-1 and erg s-1.
Put together, the temperatures, kinematics, and energetics of the outflows can provide clues to their origin. As shown in Figure 6, we find the highest warm gas fractions close to the S nucleus, with fractions decreasing at larger radii, an indication of heating from the nuclear source. Moreover, in addition to the symmetry to the north and south of the S nucleus, the fast velocities (103 km s-1), large mass loading factor (), and power ( erg s-1) found here are all typical of AGN-driven outflows (Rodríguez-Ardila et al., 2006; Müller-Sánchez et al., 2011; Hill & Zakamska, 2014; Bae & Woo, 2016; Harrison et al., 2016; Veilleux et al., 2020).
@@ 5.4.2 Feedback on Star Formation
It is well known that sustained outflows can suppress star formation (e.g., King & Pounds, 2015; Armus et al., 2020). A commonly used IR tracer of star formation are PAHs (Peeters et al., 2004; Inami et al., 2018; Lai et al., 2020, 2022). Because PAH grains are fragile, they can easily be destroyed in harsh radiation environments. As such, we can utilize PAH fluxes and ratios to assess if the outflows have any significant feedback on the local ISM.
@@
In the right panel of Figure 7, we plot 3.3 m PAH flux contours over the excitation map. Most of the PAH flux is centered around the S nucleus with emission extending to the east and west, likely tracing the disk of the southern galaxy. The emission seems to decrease in the outflow regions but the limited FOV makes it difficult to ascertain the exact correlation between the outflow and PAH emission. As such, we can take advantage of a larger FOV of MIRI to examine the PAH at 6.2 m, as shown in the flux map of Figure 8.
Like PAH 3.3 m, we measure the strongest PAH emission at the nuclear region of each galaxy, with other prominent emission seen in the disks and spiral arms, likely caused by the star formation occurring there. As was done in Figure 4, we overlay the outflowing gas as traced by the H2 S(1) 17.0 m. Compared to the regions outside the outflows and nuclei, the median PAH flux in the regions coincident with the redshifted outflow is similar: log(/W m-2) = -17.5 compared to -17.4 (see upper-right panel of Figure 8). Fluxes within the blue outflow, however, have a lower median value: log(/W m-2) = -17.9. Looking at the larger-scale structure of this merging system, we note that the northern galactic disk is roughly face on, , while the southern disk is edge on, (Sakamoto et al., 2014). Based on this geometry and the location of the nuclei, the lower PAH fluxes within the blue outflow are likely due to the outflow not being co-spatial nor in projection with the inner, more dense regions of the northern disk or spiral arm structure where PAH emission is enhanced due to active star formation. The red outflow, on the other hand, is in projection and potentially co-spatial with these inner regions of the northern galaxy.
The PAH grain size and ionization state, which can be traced through the PAH6.2/PAH7.7 and PAH11.3/PAH7.7 ratios, can also provide insight to the state of the ISM and local star formation (Rigopoulou et al., 2021; Lai et al., 2022). To calculate these ratios in NGC3256, we use the CAFE Region Extraction Tool Automaton (CRETA333https://github.com/GOALS-survey/CAFE/tree/master/CRETA, Diaz-Santos et al., in prep.) to perform grid extractions covering the full FoV of MIRI channel 1 (see Figure 9 in the Appendix). In order to robustly fit the 6.2 m, 7.7 m complex, and 11.3 m complex PAH emission features with the continuum, we opted to use the Continuum And Feature Extraction (CAFE444https://github.com/GOALS-survey/CAFE, Diaz-Santos et al., in prep.) code. PAH flux ratios were calculated based on these fits and we plot these in the spatial maps of Figure 8. These ratios are also listed in Table 2 of the Appendix.
In the lower-right panel of Figure 8, we plot PAH11.3/PAH7.7 versus PAH6.2/PAH7.7. Model tracks of theoretical values for neutral and ionized PAHs from Draine et al. (2021) are also overplotted, where we have assumed the Bruzual & Charlot (2003) interstellar radiation field with a radiation strength parameter logU = 2. Grids containing outflow emission have also been highlighted. The range of these ratios in the outflow regions (mean values of 0.21 and 0.54 for PAH6.2/PAH7.7 and PAH11.3/PAH7.7, respectively) are all within those of the disk values: 0.12 – 0.27 and 0.30 – 0.85. As such, although the outflows we detect are fast and significantly heat the molecular gas, they show no clear signs of altering the average PAH ionization state or size compared to other regions in the disk of NGC 3256, nor do they significantly reduce the surface brightness of the PAH emission in the regions of the wind. However, it is difficult to ascertain whether this lack of feedback is due to the outflow crossing over the galaxy disk (in projection) or if it is co-spatial with the disk but unable to alter the PAH grains.
@@ 6 Conclusion
Utilizing the exceptional resolution and sensitivity of JWST, we investigate the nearby merging LIRG NGC 3256 on scales of 100 pc. Here, we showcase the capability of JWST to spatially and spectrally resolve the MIR H2 lines, allowing us to directly identify and trace the outflowing warm H2 gas. Using the H2 lines from 5.0–17.0 m, we analyze the mass, temperature distribution, kinematics, and energetics of the outflowing warm H2 gas, and assess the impact of the outflows on the local star formation. The following is a summary of our findings.
• Warm, outflowing H2 gas is detected in a collimated outflow originating from the southern nucleus. The outflows extend out to a distance of 700 pc, and have a deprojected maximum velocity of 1,000 km s-1. The emission is most intense near the far edges of the outflows, possibly due to impacting the disk of the northern galaxy. We do not detect any significant outflowing H2 gas originating from the northern nucleus.
• Using the extensive wavelength coverage of JWST, we use the full set of S(8)–S(1) H2 lines to directly calculate the outflowing warm H2 gas mass: = (1.40.2)10, where the / mass fraction is 6. With an outflow time scale of about yr, the outflow mass rate is about = 2.00.8 M⊙ yr-1. These masses yield an outflow energy of erg and kinetic power of erg s-1. The mass loading factor of the warm outflowing gas is roughly unity. However, when taking into account the cold H2 component, it could be five or higher.
• The spatial map of the outflowing H2 S(7)/S(1) flux ratio reveals higher ratios closer to the S nucleus. This indicates a larger fraction of warmer gas is closer to the nucleus and decreases as the outflow expands outwards, suggesting that the S nucleus is the heating source of the outflowing gas.
• Analysis of the gas excitation in the outflow regions using [Fe II]/Pa and H2/Br line ratios shows enhanced ratios compared to those seen in typical star forming regions, indicating the outflow is energizing these regions, possibly as a shock. The value of the power-law index is also indicative of shocks, particularly in the blueshifted outflow.
• Based on 6.2 m PAH fluxes, we see no clear signs of the outflows significantly reducing the surface brightness of the PAH emission. Furthermore, analysis of the 6.2/7.7, and 11.3/7.7 PAH ratios do not show significant differences in the PAH ionization state and grain size between the outflow regions and those elsewhere in the disk.
@@
We thank the referee for their insightful feedback that helped improve the manuscript. The JWST data presented in this paper were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program 1328 and are supported by NASA grant ERS-01328. They can be accessed at 10.17909/8mkr-xr82 (catalog https://doi.org/10.17909/8mkr-xr82). STScI is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS5–26555. Support to MAST for these data is provided by the NASA Office of Space Science via grant NAG5–7584 and by other grants and contracts. T.B. and H.I. acknowledge support from JSPS KAKENHI Grant Number JP21H01129 and the Ito Foundation for Promotion of Science. A.S.E. and S.T.L. acknowledge support from NASA grants HST-GO15472 and HST-GO16914. V.U. acknowledges funding support from NSF Astronomy and Astrophysics Grant (AAG) No. AST-2408820, NASA Astrophysics Data Analysis Program (ADAP) grant No. 80NSSC23K0750, and STScI grant Nos. HST-AR-17063.005-A, HST-GO-17285.001-A, and JWST-GO-01717.001-A. C.R. acknowledges support from Fondecyt Regular grant 1230345 and ANID BASAL project FB210003 The Flatiron Institute is supported by the Simons Foundation. S.A. gratefully acknowledges support from ERC Advanced Grant 789410, the Swedish Research Council, and the Knut and Alice Wallenberg (KAW) foundation. F.M-S. acknowledges support from NASA through ADAP award 80NSSC19K1096. This paper makes use of the following ALMA data: ADS/JAO.ALMA2011.0.00525.S and 2015.1.00902.S. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), NSTC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ.
The following section presents the CRETA grid extractions as well as the PAH6.2/PAH7.7 and PAH11.3/PAH7.7 ratio maps, as discussed in Section 5.4.2. The grid sizes used to calculate the PAH ratios are , consistent with the PSF size at 11 m, and are shown in Figure 9. Two sets of grids were used, each centered on the nuclei, and they cover the full MIRI channel FoV. H2 S(1) 17.0 m outflow flux contours are overlaid to provide a sense of scale of each grid and help identify the ratios within the outflow regions. These ratios, along with those of the nuclei and disks, are listed in Table 2. The mean PAH6.2/PAH7.7 and PAH11.3/PAH7.7 ratios of the outflow regions (0.21 and 0.54) are within the ranges of the disk regions (0.12 – 0.27 and 0.30 – 0.85), suggesting the outflows are not significantly altering the average PAH grain size nor their ionization state.
| Region | PAH6.2/PAH7.7 | PAH11.3/PAH7.7 |
| Northern Nucleus | 0.270.02 | 0.300.01 |
| Southern Nucleus | 0.310.08 | 0.210.08 |
| Red Outflow N | 0.170.03 | 0.630.03 |
| 0.230.05 | 0.370.05 | |
| Red Outflow S | 0.220.04 | 0.490.02 |
| 0.210.02 | 0.620.02 | |
| 0.180.05 | 0.540.05 | |
| 0.250.02 | 0.530.03 | |
| Blue Outflow | 0.210.02 | 0.570.02 |
| Disk Regions | 0.12 – 0.27 | 0.30 – 0.85 |
Note. — Columns: (1) Locations of the extracted regions. (2-3) PAH6.2/PAH7.7 and PAH11.3/PAH7.7 flux ratios within each grid. For the disk regions, we list the range of ratio values.